OpenStudy (anonymous):

To open these doors, you must speak three functions in standard form. One function, f(x), with two real rational solutions. One function, g(x), with two real irrational solutions. One function, h(x), with two complex solutions.

OpenStudy (anonymous):

it is Completing the square

OpenStudy (anonymous):

Weird doors. But anyway, are those the only requirements?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

Ok, well, what are the two possible values for \(\sqrt{4}\)?

OpenStudy (anonymous):

2 and -2?

OpenStudy (anonymous):

Yes. So an equation like \(x^2 = 4\) would satisfy \(f(x)\). Or since it has to equal 0 to solve it, \(f(x) = x^2 - 4\)

OpenStudy (anonymous):

ohh! okay:) thankss

OpenStudy (anonymous):

You probably get the method and can do the other two now. :D

OpenStudy (anonymous):

yes for sure!:) thanks mann

OpenStudy (anonymous):

@SACAPUNTAS , hey one more question.. what is an irrational solution?

OpenStudy (anonymous):

A solution where the number is irrational. Do you know what an irrational number is?

OpenStudy (anonymous):

no:(( sorry

OpenStudy (anonymous):

I don't know the exact formal definition. But it's basically any number you can't write as a nice fraction or decimal.

OpenStudy (anonymous):

okay so would it be okay if you explained the second equation to me? i just wanna know so i can have a good grade in math, ya know?

OpenStudy (anonymous):

Well, this one is actually rather easy. Think about it! Most square roots are hideously ugly numbers.

OpenStudy (anonymous):

For example, the solutions to \(x^2 =2\) cannot be written out in any way that looks nice.

OpenStudy (anonymous):

(Except, of course, as \(x = \pm\sqrt{2}\) but you see what I mean.

OpenStudy (anonymous):

yeah! so an equation i could use would be like g(x)=?

OpenStudy (anonymous):

Look at \(f(x)\) up there and you tell me! It's the same form. :D

OpenStudy (anonymous):

g(x)= 2x^2?

OpenStudy (anonymous):

Why did you multiply. Up there you subtracted.

OpenStudy (anonymous):

idk, i am so confused.

OpenStudy (anonymous):

Ok. Want me to try explaining it again?

OpenStudy (anonymous):

i get x^2=0, but to have it setup with g(x)

OpenStudy (anonymous):

x^2=2 i ment..

OpenStudy (anonymous):

That's the part I'll explain!

OpenStudy (anonymous):

then yes explain away!:)

OpenStudy (anonymous):

Ok, the solutions to an equation are the points where y (or f(x) or g(x) or whatever = 0)

OpenStudy (anonymous):

There are sophisticated mathematical reasons for this but I can't articulate them so let's just say that's the case and go with it.

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

In other words, if we want to stick f(x) or g(x) or whatever in there, we have to replace a 0.

OpenStudy (anonymous):

How can we make one side 0 when \(x^2 = 2\)?

OpenStudy (anonymous):

subtract 2!!:) your a genius!! okay so g(x)=x^2-2?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

i wish i could give you a nether metal!

OpenStudy (anonymous):

I wish you could too. :P

OpenStudy (anonymous):

:) hah

OpenStudy (anonymous):

Anyway... do you get complex numbers?

OpenStudy (anonymous):

kind of... but can i just go over the next 2 to make sure that i got it?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

1. f(x)=x^2-4 2. g(x)=x^2-2 3. wait jk i dont get that. wow i feel stupid

OpenStudy (anonymous):

I'm happy to help. I don't want to just give you the answers though. (You felt better when you figured it out yourself, didn't you? :D )

OpenStudy (anonymous):

yes for sure!! just give me little crumbss

OpenStudy (anonymous):

Ok, so you get f(x) and g(x), right?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Ok, so a complex number means a number that can be written in form \(a + bi\) where i is the imaginary number.

OpenStudy (anonymous):

okay so i could just plug in any number for a and b?

OpenStudy (anonymous):

example. 2+4i=h(x)

OpenStudy (anonymous):

That's certainly valid, but I have a feeling they want something in terms of x.

OpenStudy (anonymous):

okay. so h(x)=2x+4i?

OpenStudy (anonymous):

That only has one complex solution.

OpenStudy (anonymous):

Do you know what the definition of i is?

OpenStudy (anonymous):

it is an imaginary number.. i= negative 1 in a square root. i^2= -1..is that what u mean?

OpenStudy (anonymous):

Right. So to get a complex number we just need an equation where we're trying to take the square root of something negative.

OpenStudy (anonymous):

okay.. So h(x)=-3x+11i?

OpenStudy (anonymous):

There are no squares in there! Like, look at the ones we did already. To solve \(x^2 - 4\) we took the square root of 4. To solve \(x^2 - 2\) we took the square root of 2.

OpenStudy (anonymous):

h(x)=-3x^2+11i?

OpenStudy (anonymous):

You know, the way the question is worded, that's actually a perfectly valid answer.

OpenStudy (anonymous):

thank gooodness! i thought i would never get it!:) thanks so much for your help

OpenStudy (anonymous):

Well, I don't think it's actually what they're going for... but it's also not wrong. So I'm a little torn on what to say. :P

OpenStudy (anonymous):

okay im going with it!!!:)

OpenStudy (anonymous):

Ok. If you get marked wrong, point out that the answer simply asked for two complex solutions, and did not specify the function had to contain only real numbers.

OpenStudy (anonymous):

okay, i really cannot thank you enough!!:D